Bell Company, a manufacturer of audio systems, started its production in October 2017. For the preceding 3 years, Bell had been a retailer of audio systems. After a thorough survey of audio system markets, Bell decided to turn its retail store into an audio equipment factory.

Raw materials cost for an audio system will total $75 per unit. Workers on the production lines are on average paid $13 per hour. An audio system usually takes 6 hours to complete. In addition, the rent on the equipment used to assemble audio systems amounts to $5,260 per month. Indirect materials cost $5 per system. A supervisor was hired to oversee production; her monthly salary is $3,780.

Factory janitorial costs are $2,060 monthly. Advertising costs for the audio system will be $9,050 per month. The factory building depreciation expense is $6,960 per year. Property taxes on the factory building will be $8,520 per year.

Assuming that Bell manufactures, on average, 1,500 audio systems per month, enter each cost item on your answer sheet, placing the dollar amount per month under the appropriate headings. Total the dollar amounts in each of the columns.

Product Costs


Cost Item

Direct
Materials

Direct
Labor

Manufacturing
Overhead

Period
Costs

Raw materials $

$

$

$

Wages for workers

Rent on equipment

Indirect materials

Factory supervisor’s salary

Janitorial costs

Advertising

Depreciation on factory building

Property taxes on factory building

$

$

$

$
Answer #1

Calculate total amount

Cost items Direct materials Direct labor Manufacturing overhead Period Costs
Raw materials 1500*75 = 112500
Wages for workers 1500*6*13 = 117000
Rent on equipment 5260
Indirect materials 1500*5 = 7500
Factory supervisor’s salary 3780
Janitorial costs 2060
Advertising 9050
Depreciation on factory building
6960/12 = 580
Property taxes on factory building 8520/12 = 710
112500 117000 19890 9050